Solve the following (Euler-Cauchy) differential equations:
(x^4D^4+6x^3D^3+9x^2D^2+3xD+1)y=(1+lnx)^2
1
Expert's answer
2021-07-23T09:50:13-0400

By the method of Euler-Cauchy , we have to solve this differential equation -


"=(x^{4}D^{4}+6x^{3}D^{3}+9x^{2}D^{2}+3xD+1)y=(1+lnx)^{2}"


Now let , "x=e^{t}" , Taking log both the side of given equation , we get ,

"=lnx=t"


Now we know that "," the terms on left hand side of given differential equation can be reduces into linear differential equation , by putting the formulas -


"=x^{4}D^{4}=D(D-1)(D-2)(D-3)y"


"=x^{3}D^{3}=" "D(D-1)(D-2)y"


"=x^{2}D^{2}=D(D-1)y"


"=xD=Dy"


Putting all these above values in the given equations , we get -



"=" "(D(D-1)(D-2)(D-3)+6D(D-1)(D-2)+9D(D-1)+3D+1)y=(1+t)^{2}"


Now , solving we get as -


"=" "(D^{4}+2D^{2}+1)y=(1+t)^{2}"


It's Auxiliary equation will be -


"=m^{4}+2m^{2}+1=0"


"=m^{4}+m^{2}+m^{2}+1=0"


"=m^{2}(m^{2}+1)+1(m^{2}+1)=0"


"=(m^{2}+1)^{2}=0"


"m={\pm}i"


so "CF=[(c_1+c_2x)cosx+(c_3+c_4x)sinx]"


Now to find "PI" of given function , we get -


"PI=\dfrac{1}{D^{4}+2D^{2}+1}" "(1+t)^{2}"


"PI=(1+D^{4}+2D^{2})^{-1}(1+t)^{2}"


Now using Binomial expansion of "(1+x)^{-1}=1-x+x^{2}+...."



Now applying this series in above equation we get ,


"PI=(1-D^{4}-2D^{2})(1+2t+t^{2})"


"PI=1+2t+t^{2}-4=t^{2}+2t-3"


Solution of given Partial differential equation can be written as -


"Y=CF+PI=" "[(c_1+c_2x)cosx+(c_3+c_4x)sinx]+(lnx)^{2}+2lnx-3"


which is required solution of given partial differential equation .